Write an equation of the line through the given point (-3, -6) with the given slope m = 3. Write the equation in slope-intercept form.
Found 2 solutions by solve_for_x, ewatrrr:
Answer by solve_for_x(190)

The slope-intercept form of the equation of a line is:

y = mx + b

where m is the slope of the line and b is the value of the y-intercept.

Since the slope is given as m = 3, the line will have the form:

y = 3x + b

To determine the value of b, substitute the known point (-3, -6) and solve for b:

-6 = 3(-3) + b

-6 = -9 + b

b = -6 + 9

b = 3

The equation of the line is then y = 3x + 3.

Answer by ewatrrr(24785)

P(-3,-6), m = 3
y + 6 = 3(x+3)
y = 3x + 3 in slope-intercept form.